OpenStudy (anonymous):

When Tim started football 3 years ago, his coach was twice his age. Currently, Tim's coach is 10 years older than him. How old is Tim now?

OpenStudy (anonymous):

tim's age = x coach = y 2(x-3) = y -3 .......1 x = y - 10 solve the equations